OpenStudy (spareb665):

The average annual number of cigarettes smoked by an adult in some country continues to decline. For the years after 1997 and y is the average annual number of cigarettes smoked. If this trend continues, find the year in which the average annual number of cigarettes smoked is zero. To do this, let y=0 and solve for x.

OpenStudy (anonymous):

Is the question complete... :D

OpenStudy (spareb665):

it appears so..

OpenStudy (anonymous):

Isn't there a rate at which it declines...

OpenStudy (spareb665):

no thats the entire problem word for word

OpenStudy (spareb665):

its really confusing

OpenStudy (anonymous):

No..you missed the question somewhere..the rate is not given..

OpenStudy (anonymous):

You should have some graph or equation for solving this.

OpenStudy (spareb665):

The average annual number of cigarettes smoked by an adult in some country continues to decline. For the years 1997-2006, the equation y=-49.6x +1756.7 approximates this data. Here, x is the number of years after 1997 and y is the average annual number of cigarettes smoked if this trend continues find the year in which the average annual number of cigarettes smoked is zero. To do this , let y=0 and solve for x.

OpenStudy (anonymous):

Put y=o 0 = -49.6x + 1756.7 49.6x = 1756.7 x = 35.4 Since, x is the number of years after 1997..add the value of x to 1997 to get the year in which y becomes zero...So, the answer is 1997 + 35.4 = 2032.4 So, the year is 2032

OpenStudy (anonymous):

@spareb665

OpenStudy (spareb665):

thanks abhi!

OpenStudy (anonymous):

welcome : )